Q.4. A robot lands on Mars. Its x and y coordinates are given by equations:

x(t) = 2m –(0.25m/s2

)t2

y(t)= (1.0 m/s)t +(0.025m/s3

)


(a).What is the position and distance from landing at t=2 seconds?(b)What is the

displacement and average velocity during the interval from t=0 s to t= 2?(c).Derive

a general expression for its instantaneous velocity vector and find its instantaneous

velocity in its component form as well as in terms of its magnitude.


1
Expert's answer
2021-10-10T16:06:09-0400

a) "x(2)=2-0.25\cdot2^2=1\ (m)"

"y(2)=1\cdot 2+0.025\cdot2^3=2.2\ (m)"


"s=\int_0^2((0.5t)^2+(1+0.075t^2)^2)^{0.5}dt=2.46\ (m)"


b) "d=\sqrt{1^2+2.2^2}=2.42 (m)"


"v_{avr}=2.42/2=1.21\ (m/s)"


c) "\vec{v}=-0.5t\cdot\vec{i}+(1+0.075t^2)\cdot\vec{j}"


"v_x=-0.5t" and "v_y=1+0.075t^2"


"|\vec{v}|=\sqrt{(-0.5t)^2+(1+0.075t^2)^2}"